OpenStudy (anonymous):

Water flowing over Lower Falls in Yellowstone National Park drops 94m. If all the water’s energy goes to heat it, what is its temperature increase?

OpenStudy (shamim):

Can u tell me the potential energy at 94m above

OpenStudy (shamim):

The formula is P.E=mgh=m*9.8*94 J

OpenStudy (shamim):

This potential energy will spend to increase the temperature of water

OpenStudy (shamim):

U know Q=ms*del theta

OpenStudy (shamim):

So m*9.8*94=ms*del theta 9.8*94=4200*del theta Del theta=?

OpenStudy (shamim):

Response plz!!!!!!!